The proportion of eligible voters in the next corporation election (in tumkur) who will vote 

for the incumbent is assumed to be 55%. What is the probability that in a random sample of 

500 voters less than 49% say they will vote for the incumbent? Write your conclusion. 


1
Expert's answer
2022-03-11T02:02:11-0500

we define the following

Assumptions/Conditions

Randomization: the 500 voters were chosen in a random sample

Independence: one person's vote does not affect another's

Sample Size: there are more than 5000 voters

Success/Failure:

np = (500)(0.55) > 10

nq = (500)(0.45) > 10


therefore we can use the normal model

mean = 0.55

standard deviation = ( (0.55*0.45)/500 )1/2= 0.0222

N(0.55, 0.0222)


z = (0.49 - 0.55) / (0 .0222) = -2.703

according to the standard z-table, the probability that z<-2.703 is 0.0035 which equals (0.35%)


Hence we may conclude that there is a 0.35% probability that in a random sample of 

500 voters less than 49% say they will vote for the incumbent.